9. The British government has a consol bond outstanding paying £200 per year forever. Assume the current interest rate is 12% per year a. What is the value of the bond immediately after a payment is made? b. What is the value of the bond immediately before a payment is made? a. What is the value of the bond immediately after a payment is made? The value of the bond immediately after a payment is made is The Value of the bond immediately after a payment is made is£ Round to the nearest pound.) b. What is the value of the bond immediately before a payment is made? The value of the bond immediately before a payment is made is (Round to the nearest pound.)
Answer #1

a)After:  PV = 200/0.12 = 1666.67

b)Before : PV = 200 + 200/0.12 = 1867.67

answered by: mai đặng